# Surrogate-based multilevel Monte Carlo methods for uncertainty quantification in the Grad-Shafranov free boundary problem

### v1

@@ 4.4 Properties of geometric parameters
Refer to caption Refer to caption Refer to caption Refer to caption
 MC-FE + Direct Solve MC-FE + Surrogate MLMC-FE + Direct Solve MLMC-FE + Surrogate
@@
Figure 4: The overlayed plasma boundaries of 50 random realizations are displayed in the top row as violet curves (interpolated to the neighboring finer mesh). The solid violet line is the plasma boundary of the expected poloidal flux generated with tolerance $\epsilon=4\times 10^{-4}$. The inner and outer walls of the reactor are displayed in solid black and dark red respectively. The bottom row shows the regions close to the x-points in more detail. The dark green dots are the x-points of the expected solution. The columns from left to right correspond to simulations using the MC-FE approach with the direct solver and surrogate, MLMC-FE with direct solver and surrogate. All simulations were performed using the discretization level $\ell=5$ on geometry-conforming uniform meshes.
@@
Plasma boundary. Figure 4 shows the plasma boundary of the expected poloidal flux in (3) for both direct solve and surrogate evaluation. As discussed in [12], Monte Carlo sampling with direct computation on a single spatial grid results in a smooth plasma boundary, and the use of a surrogate achieves the same outcome, as seen in the second column of the figure. However, multilevel Monte Carlo sampling on non-nested geometry-conforming meshes introduces significant distortion near the x-point due to extrapolation errors arising during the accumulation of sample corrections, evident in the third and fourth columns of Figure 4. This distortion can be eliminated by interpolating the multilevel solutions to a common fine grid (near the x-point, solutions computed this way look exactly like those in the bottom-left two images of Figure 4), but as shown in Table 4, this comes at significant expense. Thus, the use of multilevel methods in this setting exhibits a trade-off between dramatically increased efficiency and somewhat decreased accuracy.
@@
Geometric descriptors. Table 6 summarizes some values of geometric parameters derived from the plasma boundaries of the expected plasma field for various methods. In the table, we set the outcome of Monte Carlo sampling with direct computations as the benchmark. We find that the surrogate-enhanced Monte Carlo produces parameters identical to the direct-solve results up to two decimal places. In contrast, MLMC sampling with both direct computation and surrogate methods on geometry-conforming grids aligns only up to one decimal place, due to plasma boundary distortions from extrapolation errors on non-nested grids. However, when using a common grid for interpolation, the geometric descriptors for MLMC sampling with both direct solve and surrogate match up to two decimal places.
MC-FE DS MC-FE Surrogate MLMC-FE DS MLMC-FE Surrogate MLMC-FE DS (Interp) MLMC-FE Surrogate (Interp)
x point (5.14,-3.29) (5.14,-3.29) (5.14,-3.29) (5.14, -3.30) (5.14,-3.29) (5.14,-3.29)
magnetic axis (6.41,0.61) (6.41,0.61) (6.44,0.56) (6.44, 0.56) (6.41,0.61) (6.41,0.61)
strike (4.16,-3.71) (4.16,-3.71) (4.16,-3.71) (4.16, -3.71) (4.16,-3.71) (4.16,-3.71)
points (5.56,-4.22) (5.56,-4.22) (5.56,-4.22) (5.56, -4.22) (5.56,-4.22) (5.56,-4.22)
inverse aspect ratio 0.32 0.32 0.32 0.32 0.32 0.32
elongation 1.86 1.86 1.87 1.87 1.86 1.86
upper triangularity 0.43 0.43 0.43 0.43 0.43 0.43
lower triangularity 0.53 0.53 0.53 0.53 0.53 0.53
Table 6: Geometric parameters of the expected poloidal flux $u$ from MC-FE with direct solver, MC-FE with the surrogate, MLMC-FE with direct solver, MLMC-FE with the surrogate, MLMC-FE with direct solver with interpolating solution to a common fine grid of level $L=5$, MLMC-FE with surrogate with interpolating solution to a common fine grid of level $L=5$. The results are generated with an nMSE $4\times 10^{-4}$ on the geometry-conforming uniform mesh set.
@@ 5 Concluding remarks
This paper proposes a surrogate-based MLMC-FE sampling strategy to estimate the expectation of the magnetic field in the Grad-Shafranov free boundary problem with high-dimensional uncertainties in current intensities. Cost analyses demonstrate the effects of using surrogate approximations defined by sparse-grid collocation methods to the solutions of the nonlinear systems of equations arising in the model, or of using multilevel Monte Carlo methods to reduce the cost of direct solves, and of combining these two ideas to use sparse-grid surrogates together with multilevel methods. Computational experiments with the Grad-Shafranov equation demonstrate that each of these approaches leads to reductions in costs to perform simulations, with very dramatic cost reductions (approaching factors of $10^{4}$) obtained from the combined methods. These savings come with some sacrifice of accuracy for the surrogate solutions, most notably for multilevel spatial methods, although the quantitative values of important measures in the model agree to two digits.

### v2

@@ 4.4 Properties of geometric parameters
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-FE MC-FE MLMC-FE MLMC-FE
Direct Solver Surrogate Direct Solver Surrogate
@@
Reactor MLMC-FE MLMC-FE MLMC-FE MLMC-FE
cross section Direct Solver Surrogate Direct Solver Surrogate
Heat flow Heat flow Interpolation Interpolation
@@
Figure 4: The overlayed plasma boundaries of 50 random realizations are displayed as violet curves. The solid violet line is the plasma boundary of the expected poloidal flux generated with tolerance $\epsilon=4\times 10^{-4}$. On the leftmost panel, the inner and outer walls of the reactor are displayed in solid black and dark red respectively, while the green square highlights the zoomed–in in the remaining images. The dark green dots are the x-points of the expected solution. All simulations were performed using the discretization level $\ell=5$ on geometry-conforming uniform meshes.
@@
Plasma boundary. Figure 4 shows the plasma boundary of the expected poloidal flux in (3) for both direct solve and surrogate evaluation. As discussed in [14], Monte Carlo sampling with direct computation on a single spatial grid results in a smooth plasma boundary, and the use of a surrogate achieves the same outcome, as seen in the second and third columns of the top row in the figure. However, multilevel Monte Carlo sampling introduces distortions in the profile of the plasma boundary due to the interaction of samples defined across different grids during the accumulation of sample corrections. This is evident in the last two columns of the top row in Figure 4. Thus, the straightforward use of multilevel methods in this setting exhibits a trade-off between dramatically increased efficiency and somewhat decreased accuracy.
Post–processing strategies. We propose two strategies to eliminate this distortion. The principled approach is to interpolate the sample corrections to a common fine grid (near the x-point, solutions computed this way look exactly like those in the bottom-right two images of Figure 4), but as shown in Table 4, this comes at significant expense. A second strategy to filter out the spurious high–frequency components of the multi–level approximation is to take advantage of the smoothing properties of the heat operator by subjecting the approximation to a simple heat flow. This approach avoids the expense of interpolating every realization into a common grid.
More precisely, let us denote the multi–level approximation to the expectation as $\psi$ and by
[…]
the mass and stiffness matrices for the nodal basis functions $\{\phi_{i}\}$ of the mesh where the approximated expectation $\psi$ was computed. Then, the filtered approximation $\psi^{*}$ is given by
$\psi^{*}=\left(\left(\mathbf{M}+\kappa\alpha\mathbf{S}\right)^{-1}\mathbf{M}\right)^{n}\psi\,.$
The matrix $\mathbf{A}:=\left(\mathbf{M}+\kappa\alpha\mathbf{S}\right)$ can be precomputed once and then the system $\mathbf{A}^{-1}\mathbf{M}\psi$ solved $n$ times to obtain the filtered expectation $\psi^{*}$. The post–processing formula above amounts to an implicit–in–time Galerkin discretization of the initial boundary value problem
$\partial_{t}\psi^{*}=\alpha\Delta\psi^{*}\;\;\text{ in }\Omega\times(0,n\kappa]\,,\qquad\partial_{\nu}\psi^{*}=0\;\;\text{ on }\Gamma\times(0,n\kappa]\,,\qquad\psi^{*}=\psi\;\;\text{ on }\overline{\Omega}\times\{0\}\,,$
where $0<\kappa$ is the size of the pseudo time step, $0<\alpha$ is the pseudo diffusivity, and $n$ is the number of pseudo time steps. In our experiments we used $\kappa=2\times 10^{-2}$, $\alpha=0.1$, and $n=5$.
As depicted in the second and third columns in the bottom row of Figure 4, after the post–processing step, the distortion of the boundary essentially disappears at a computational cost that is significantly smaller than that of interpolating the samples into a common grid. As shown in Table 4, the associated speed–ups (final column) are not quite as large as those of the un–processed strategy (seventh column) but still significant compared to those of the multi–level direct solver (fourth column). This is specially true for the more challenging cases of stringent tolerances, where the computational time was decreased by factors ranging between $10^{2}$ and $10^{4}$. As we will discuss in the next subsection, the post–processing step has no detrimental effect on the determination of the relevant geometric parameters of the plasma.
@@
Geometric descriptors. Table 6 summarizes some values of geometric parameters derived from the plasma boundaries of the expected plasma field for various methods. In the table, we set the outcome of Monte Carlo sampling with direct computations as the benchmark. We find that the surrogate–enhanced Monte Carlo produces parameters identical to the direct–solve results up to two decimal places. In contrast, MLMC sampling with both direct computation and surrogate methods on geometry–conforming grids aligns only up to one decimal place, due to plasma boundary distortions from extrapolation errors on non–nested grids. However, when using a common grid for interpolation, the geometric descriptors for MLMC sampling, both direct solve and surrogate, incur a relative error significantly smaller than 1%. A similarly accurate behavior is observed with both filtering strategies for the multi–level approximations.
Interpolation Interpolation Heat flow Heat flow
MC-FE MC-FE MLMC-FE MLMC-FE MLMC-FE MLMC-FE MLMC-FE MLMC-FE
Direct Solver Surrogate Direct solver Surrogate Direct solver Surrogate Direct solver Surrogate
x point (5.14,-3.29) (5.14,-3.29) (5.14,-3.29) (5.14, -3.30) (5.14,-3.29) (5.14,-3.29) (5.14,-3.28) (5.14,-3.28)
magnetic axis (6.41,0.61) (6.41,0.61) (6.44,0.56) (6.44, 0.56) (6.41,0.61) (6.41,0.61) (6.41,0.60) (6.41,0.60)
strike (4.16,-3.71) (4.16,-3.71) (4.16,-3.71) (4.16, -3.71) (4.16,-3.71) (4.16,-3.71) (4.16,-3.71) (4.16,-3.71)
points (5.56,-4.22) (5.56,-4.22) (5.56,-4.22) (5.56, -4.22) (5.56,-4.22) (5.56,-4.22) (5.56,-4.22) (5.56,-4.22)
inverse aspect ratio 0.32 0.32 0.32 0.32 0.32 0.32 0.32 0.32
elongation 1.86 1.86 1.87 1.87 1.86 1.86 1.86 1.86
upper triangularity 0.43 0.43 0.43 0.43 0.43 0.43 0.44 0.44
lower triangularity 0.53 0.53 0.53 0.53 0.53 0.53 0.53 0.53
Table 6: Geometric parameters of the expected poloidal flux $u$. The results are generated with a target normalized mean squared error of $\epsilon=4\times 10^{-4}$. In the case of the interpolation post–processing strategy, the results were interpolated into a common fine grid of level $L=5$.
@@ 5 Concluding remarks
This paper proposes a surrogate-based MLMC-FE sampling strategy to estimate the expectation of the magnetic field in the Grad-Shafranov free boundary problem with high-dimensional uncertainties in current intensities. Cost analyses demonstrate the effects of using surrogate approximations defined by sparse-grid collocation methods to the solutions of the nonlinear systems of equations arising in the model, or of using multilevel Monte Carlo methods to reduce the cost of direct solves, and of combining these two ideas to use sparse-grid surrogates together with multilevel methods. Computational experiments with the Grad-Shafranov equation demonstrate that each of these approaches leads to reductions in costs to perform simulations, with very dramatic cost reductions (approaching factors of $10^{4}$) obtained from the combined methods. These savings come with some sacrifice of accuracy for the surrogate solutions, most notably for multilevel spatial methods, although the quantitative values of important measures in the model agree to two digits. Two filtering strategies are offered to counteract the loss of accuracy in the multi–level estimations. The heat flow smoothing is by far the the most efficient, as it offers speed–ups ranging from $10^{2}$ to $10^{4}$ and produces estimates that fall within less of 1% of those of a direct Monte Carlo estimation on the finest spatial grid.